happ company had $230,000 in assets and $70,000 in liabilities on january 1, 2019. During 2019 Happ's assets decrease $20,000 and its loabilities increase $10,000. What are the balances for Happs assets,liabilities, and stockholders equity on december 31,2019?
Answer #1

Total assets on December 31,2019=(230,000-20,000)=$210,000

Total liabilities on December 31,2019=(70,000+10,000)=$80,000

Total assets=Total liabilities+Total equity

Total equity on December 31,2019=(210,000-80,000)=$130,000